# FiVE [Uncaptioned image]: A Fine-grained Video Editing Benchmark for Evaluating Emerging Diffusion and Rectified Flow Models

### v1

@@ 1 Introduction
Table 1: Statistics of existing video editing datasets. ‘DV’ and ‘YT’ represent and DAVIS dateset and YouTube website, respectively.
Dataset Num. Source Gen- Num. Src. obj. Edit obj. Obj. Src. obj. Type of Edited Prompts
Videos of Videos Video Prompts Words Words Instruct. Masks
TGVE [48] 76 DV, YT, Videvo 304 style, object, background, multiple
TGVE+ [34] 76 DV, YT, Videvo 1417 style, object, background, multiple, color, add, remove
FiVE (Ours) 100 DV, GenVideo 420 object (w&w/o non-rigid), color, material, add, remove
@@ 2.1 Diffusion and RF Inversion
Diffusion inversion. Diffusion models (DMs) drive advances in inversion techniques for better control over generated samples. DDIM [35] accelerates sampling but struggles with fine-grained reconstructions due to nonlinearities and score estimation errors. To improve accuracy, fine-tuning and optimization-based methods [25, 7, 40, 46, 16] have been proposed. While these advancements significantly enhance the fidelity and accuracy of DDIM inversion, the associated computational costs and resource demands remain substantial limiting factors. Although diffusion-based methods have achieved outstanding performance, recent studies [8, 32, 20, 15] suggest that Rectified Flow models (RFs) hold significant potential to surpass DMs in certain applications. Notably, the inversion and editing capabilities of RF models have been explored in image editing [32, 19, 5, 42], demonstrating their efficiency and effectiveness. Building on these strengths, this work extend their potential in video editing.
@@ 2.2 Video Editing
Massive data fine-tuned models undergo extensive training on large datasets for high-quality, versatile video editing [29]. InstructVid2Vid [29] enables complex edits via natural language instructions, while EffiVED [54] refines broad datasets into high-quality subsets for efficient editing. SF-V [55] introduces adversarial training for video synthesis. Despite their precision and adaptability, these methods are computationally expensive, limiting practicality in resource-constrained settings. These methods focus on learning video motion, which may compromise background preservation. Recently, VideoGrain [50] has emphasized multi-granularity video editing rather than fine-grained edits, aiming to retain background details while modifying target objects. This work aims to address this issue and advance the field.
@@ 3.1 FiVE Benchmark
Six Fine-grained Video Editing Tasks. FiVE comprises six fine-grained video editing tasks, totaling 420 high-quality editing prompt pairs. It primarily includes four object-targeted editing tasks of increasing complexity: color alteration, material modification, object substitution without non-rigid deformation, and object substitution with non-rigid deformation. For each task, we use GPT-4o [13] to generate four source-target prompt pairs per video in the benchmark, resulting in 400 high-quality editing pairs. To further enhance the diversity of editing types within the benchmark, we select 10 videos each for the add and remove tasks and design corresponding source-target prompt pairs. Additionally, we explicitly provide both object-related source words and edited words in the source and edited text prompts, respectively, denoted as ‘src. obj. words’ and ‘edit obj. words’ in Table 1. To accommodate different editing methods, we use GPT-4o [13] to generate instruction prompts (‘Obj. Instr.’ in Table 1) for each source-target pair, ensuring compatibility with models like InstructPix2Pix [2]. For each video, we employ SAM2 [30] to generate masks of the edited regions, enabling the evaluation of background preservation metrics. Additional details can be found in the Appendix.
@@ 4.1 Preliminary
Here, the vector field $v_{t}(X_{t})=f(X_{t},c,1-t;\varphi)$, where $f(\cdot)$ is a neural network parameterized by $\varphi$ and $c$ is the guided text. Given the ODE in Eq. (1), we start with $X_{0}\sim q_{0}$ from the source distribution and integrate over $t:0\rightarrow 1$ to reach the endpoint $X_{1}\sim p_{0}$, which is sampled from the target distribution. the forward process of RF assumes the linear path between the two states: $X_{t}=(1-t)X_{0}+tX_{1}$, then we can derive the vector field of $v_{t}(\cdot)$: $u_{t}(X_{t})=X_{1}-X_{0}$. Then, $v_{t}(X_{t})$ is used to approximate the network $f(X_{t},c,t;\varphi)$ by minimizing the loss function $\mathcal{L}_{\varphi}=\small\mathbb{E}_{t\sim\mathcal{U}[0,1],Y_{t}}[\|v_{t}(X_{t})-f(X_{t},c,t;\varphi)\|_{2}^{2}].$ RF ensures that their sampling paths are relatively straight, enabling the use of a small number of discretization steps.
@@
$\triangleq X_{t-1}^{edit}+\mathbb{E}[v^{\Delta}_{t}(\hat{X}_{t}^{tgt},X_{t}^{src})|X_{1}^{src})$ (4)
@@ 4.2.2 Wan-Edit
where $\hat{\mathbf{X}}_{t}^{tgt}=\mathbf{X}_{t-1}^{edit}\!+\!\mathbf{X}_{t}^{src}\!-\!\mathbf{X}_{1}^{src}$. This allows for smooth and coherent edits across all frames, preserving the temporal dynamics of the video.
@@ 6 Conclusion
  • A. More details on the baseline methods
  • B. More details on implementation details
  • C. More details on FiVE Dataset
  • D. GPU memory and speed comparison
  • E. More quantitative results
  • F. More qualitative results
@@ E More Quantitative Results and Analysis
Tables 5 and 8 present the results for Edit1: Object replacement without non-rigid transformations (e.g., replacing a car with a bus). Our proposed Wan-Edit achieves background preservation and text alignment comparable to VideoGrain, while exhibiting superior motion fidelity. Since Wan-Edit better retains the original video background, its IQA scores remain consistent with those of the source video. Regarding the FiVE-Acc metrics, which assess the accuracy of successful edits, the most competitive method is DMT, which optimizes editing based on the input text. The training-free Wan-Edit and VideoGrain achieve similar results, slightly trailing DMT but significantly outperforming other methods. Overall, DMT offers the best text-vision alignment but requires optimization, whereas VideoGrain and Wan-Edit strike a strong balance across various fine-grained video editing metrics. Notably, Wan-Edit stands out for its superior efficiency, delivering faster and more stable results compared to VideoGrain.
@@
Tables 6 and 8 present the results of all compared methods on Edit2: Object Replacement with Non-Rigid Transformations, revealing similar conclusions to Edit1: Object Replacement with Rigid Transformations. However, Edit2 is more challenging than Edit1 due to the complexity of non-rigid transformations. This increased difficulty results in a noticeable drop in text-vision alignment, motion fidelity scores, and the editing success rate (FiVE-Acc) metrics compared to Edit1. In terms of the editing success rate (FiVE-Acc), DMT and Wan-Edit achieve 67.86% and 52.02%, respectively, on Edit1, with DMT outperforming Wan-Edit by approximately 15%. However, on Edit2, DMT drops significantly to 53.72%, while Wan-Edit remains stable. This indicates Wan-Edit’s robustness in handling non-rigid transformations, maintaining consistent performance even in more challenging editing scenarios.
@@
Similarly, consistent trends are observed across other editing types, Edit3 (color changes) and Edit4 (object material changes) as shown in Tables 9 - 12, further reinforcing our conclusions. Regarding the FiVE-Acc metric, color changes ( Edit3) achieve the highest editing success rate among all types, with VideoGrain reaching 86% and Wan-Edit at 63%. In contrast, object material changes ( Edit4) are significantly more challenging, with AnyV2V achieving the highest success rate at 43%, followed by Pyramid-Edit at 36%. This difficulty arises because object material changes often require modifying mid- and low-frequency noise, making training-free methods highly sensitive to parameters, which leads to lower success rates.
@@
Tables 13 - 16 compare the results of Edit5 (object addition) and Edit6 (object removal). For object addition ( Edit5), VideoGrain and Wan-Edit achieve the highest scores in background preservation and motion fidelity, while TokenFlow performs best in text-vision alignment and IQA. In terms of FiVE-Acc, the RF-based methods Pyramid-Edit and Wan-Edit achieve success rates of 83% and 72%, respectively, whereas the highest-performing diffusion-based method, DMT, reaches only 61%, highlighting the advantage of RF-based approaches in this task. For Edit6 (object removal), nearly all methods perform the worst among all editing types, with FiVE-Acc scores dropping below 20%, indicating that object removal is one of the most challenging fine-grained editing tasks. This difficulty arises because removing an object requires precisely inpainting the occluded background while maintaining temporal coherence, which is particularly challenging for existing editing models. The visualizations in Fig. 4 of the main paper further confirm these findings.
@@ F More Qualitative Results and Analysis
We present the editing results across various editing types and comparison methods in Figs. 10 - 12 and Fig. 13. Figs. 10 - 12 compare the results of diffusion-based and RF-based editing methods across six editing types. These comparisons highlight the strengths and weaknesses of each method while also demonstrating the superiority of VideoGrain and our Wan-Edit. The videos shown in Fig. 13 are generated examples and represent a particularly challenging editing case. Editing becomes difficult when the object occupies a significant portion of the video frame, as it requires modifying low-frequency noise while maintaining spatial and temporal consistency. This results in failure even for the best-performing Wan-Edit. To better showcase the dynamic consistency in video editing, more video demos are available on the anonymous website: https://sites.google.com/view/five-benchmark.

### v2

@@ 1 Introduction
Table 1: Statistics of existing video editing datasets and benchmarks.
Dataset Usage Num. Num. Frames Num. Gen. Src. obj. Edit obj. Obj. Src. obj. Type of Edited Prompts
Images Videos Per Video Prompts Videos Words Words Instruct. Masks
VIVID-10M Train 672K 73.7K 30 10M object, add, remove
Señorita-2M Train 0 388.9K 33$\sim$64 2M style, object, multiple, color, add, remove, motion
TGVE Eval 0 76 32 304 style, object, bg, multiple
TGVE+ Eval 0 76 32 1417 style, object, bg, multiple, color, add, remove
VIVID-10M Eval 0 64 33$\sim$64 852 object, add, remove
FiVE (Ours) Eval 0 100 35 $\sim$ 126 420 object (w&w/o non-rigid), color, material, add, remove
@@ 2.1 Diffusion and RF Inversion
Diffusion inversion. Diffusion models (DMs) drive advances in inversion techniques for better control over generated samples. DDIM [36] accelerates sampling but struggles with fine-grained reconstructions due to nonlinearities and score estimation errors. To improve accuracy, fine-tuning and optimization-based methods [26, 7, 42, 48, 16] have been proposed. While these advancements significantly enhance the fidelity and accuracy of DDIM inversion, the associated computational costs and resource demands remain substantial limiting factors. Although diffusion-based methods have achieved outstanding performance, recent studies [8, 33, 20, 15] suggest that Rectified Flow models (RFs) hold significant potential to surpass DMs in certain applications. Notably, the inversion and editing capabilities of RF models have been explored in image editing [33, 19, 5, 44, 51], demonstrating their efficiency and effectiveness. Building on these strengths, this work extend their potential in video editing.
@@ 2.2 Video Editing
Massive data fine-tuned models undergo extensive training on large datasets for high-quality, versatile video editing [30, 14, 38]. InstructVid2Vid [30] enables complex edits via natural language instructions, while EffiVED [58] refines broad datasets into high-quality subsets for efficient editing. SF-V [59] introduces adversarial training for video synthesis. Despite their precision and adaptability, these methods are computationally expensive, limiting practicality in resource-constrained settings. These methods focus on learning video motion, which may compromise background preservation. Recently, VideoGrain [53] has emphasized multi-granularity video editing rather than fine-grained edits, aiming to retain background details while modifying target objects. This work aims to address this issue and advance the field.
Recent unified architectures [14, 38, 55] have also incorporated video editing into their multi-functional frameworks. VACE  [14] enables users to perform video generation, editing, and personalization in an integrated manner. Such architectures typically concatenate reference and target tokens as input to the model, allowing for a unified processing pipeline.
@@ 3.1 FiVE Benchmark
Six Fine-grained Video Editing Tasks. FiVE comprises six fine-grained video editing tasks, totaling 420 high-quality editing prompt pairs. It primarily includes four object-targeted editing tasks of increasing complexity: color alteration, material modification, object substitution without non-rigid deformation, and object substitution with non-rigid deformation. For each task, we use GPT-4o [13] to generate four source-target prompt pairs per video in the benchmark, resulting in 400 high-quality editing pairs. To further enhance the diversity of editing types within the benchmark, we select 10 videos each for the add and remove tasks and design corresponding source-target prompt pairs. Additionally, we explicitly provide both object-related source words and edited words in the source and edited text prompts, respectively, denoted as ‘src. obj. words’ and ‘edit obj. words’ in Table 1. To accommodate different editing methods, we use GPT-4o [13] to generate instruction prompts (‘Obj. Instr.’ in Table 1) for each source-target pair, ensuring compatibility with models like InstructPix2Pix [2]. For each video, we employ SAM2 [31, 21] to generate masks of the edited regions, enabling the evaluation of background preservation metrics. Additional details can be found in the Appendix.
@@ 4.1 Preliminary
Here, the vector field $v_{t}(X_{t})=f(X_{t},c,1-t;\varphi)$, where $f(\cdot)$ is a neural network parameterized by $\varphi$ and $c$ is the guided text. Given the ODE in Eq. (1), we start with $X_{0}\sim q_{0}$ from the source distribution and integrate over $t:0\rightarrow 1$ to reach the endpoint $X_{1}\sim p_{0}$, which is sampled from the target distribution. the forward process of RF assumes the linear path between the two states: $X_{t}=(1-t)X_{0}+tX_{1}$, then we can derive the vector field of $v_{t}(\cdot)$: $u_{t}(X_{t})=X_{1}-X_{0}$. Then, $v_{t}(X_{t})$ is used to approximate the network $f(X_{t},c,t;\varphi)$ by minimizing the loss function $\mathcal{L}_{\varphi}=\mathbb{E}_{t\sim\mathcal{U}[0,1],Y_{t}}[\|v_{t}(X_{t})-f(X_{t},c,t;\varphi)\|_{2}^{2}].$ RF ensures that their sampling paths are relatively straight, enabling the use of a small number of discretization steps.
@@
$\triangleq X_{t-1}^{edit}+\mathbb{E}[v^{\Delta}_{t}(\hat{X}_{t}^{tgt},X_{t}^{src})|X_{1}^{src}]dt$ (4)
@@ 4.2.2 Wan-Edit
where $\hat{\mathbf{X}}_{t}^{tgt}\approx\mathbf{X}_{t-1}^{edit}\!+\!\mathbf{X}_{t}^{src}\!-\!\mathbf{X}_{1}^{src}$. This allows for smooth and coherent edits across all frames, preserving the temporal dynamics of the video.
@@ 5.2.1 Comparison on Common Metrics
Figure 3: Human evaluation example using Netlify. Left: An example illustrating human verification of FiVE-Acc metric, conducted on WanEdit results. Central: A human preference study where annotators select the top-2 preferred results. Right: preference statistics.
Figure 4: Human validation of mask quality.
@@ 5.2.3 Comparison on All Metrics
5.3 Human Validation
Table 4: Human vs. VLM model evaluation on FiVE-Acc metrics.
[…]
Qwen-2.5-VL 41.41 52.53 55.72 38.22 46.97
Human 44.37 50.31 51.98 42.70 47.34
To ensure the reliability and perceptual alignment of our evaluation protocol, we conduct comprehensive human validation on the proposed FiVE-Acc metric, the compared editing methods, and the quality of segmentation masks.
Human validation of FiVE-Acc metrics. In Fig. 3 (left), we conducted a human study on 16 randomly sampled videos with 64 prompts. Table 4 shows that human scores closely match Qwen-2.5-VL’s results (47.34 vs. 46.97), confirming FiVE-Acc aligns well with human perception and offers consistent evaluation.
Human validation across compared methods. In Fig. 3 (central), we randomly sample 11 videos, resulting in 45 source–target prompt pairs. All seven methods from main Table 2 (methods a–g) are evaluated. To reduce bias, the order of methods is randomized. Annotators are asked to select their top-2 preferred results per prompt, and the aggregated votes are used to compute preference statistics in Fig. 3 (right). The results indicate that Method g (our Wan-Edit) significantly outperforms the others, receiving 66.2% top-1 votes and 41.9% top-1&2 combined votes.
Human validation of mask quality. The segmentation masks in FiVE are initially generated by the SAM model to provide object-level supervision. To ensure accuracy and alignment with the intended target regions, each mask is manually reviewed and corrected if necessary by human annotators (see Fig. 4). This process guarantees high-quality annotations that support reliable evaluation and training.
@@ 6 Conclusion
  • More details on the baseline methods
  • More details on implementation details
  • More details on FiVE Dataset
  • GPU memory and speed comparison
  • More quantitative results and analysis
  • More qualitative results and analysis
@@ E More Quantitative Results and Analysis
Tables 6 and 9 present the results for Edit1: Object replacement without non-rigid transformations (e.g., replacing a car with a bus). Our proposed Wan-Edit achieves background preservation and text alignment comparable to VideoGrain, while exhibiting superior motion fidelity. Since Wan-Edit better retains the original video background, its IQA scores remain consistent with those of the source video. Regarding the FiVE-Acc metrics, which assess the accuracy of successful edits, the most competitive method is DMT, which optimizes editing based on the input text. The training-free Wan-Edit and VideoGrain achieve similar results, slightly trailing DMT but significantly outperforming other methods. Overall, DMT offers the best text-vision alignment but requires optimization, whereas VideoGrain and Wan-Edit strike a strong balance across various fine-grained video editing metrics. Notably, Wan-Edit stands out for its superior efficiency, delivering faster and more stable results compared to VideoGrain.
@@
Tables 7 and 9 present the results of all compared methods on Edit2: Object Replacement with Non-Rigid Transformations, revealing similar conclusions to Edit1: Object Replacement with Rigid Transformations. However, Edit2 is more challenging than Edit1 due to the complexity of non-rigid transformations. This increased difficulty results in a noticeable drop in text-vision alignment, motion fidelity scores, and the editing success rate (FiVE-Acc) metrics compared to Edit1. In terms of the editing success rate (FiVE-Acc), DMT and Wan-Edit achieve 67.86% and 52.02%, respectively, on Edit1, with DMT outperforming Wan-Edit by approximately 15%. However, on Edit2, DMT drops significantly to 53.72%, while Wan-Edit remains stable. This indicates Wan-Edit’s robustness in handling non-rigid transformations, maintaining consistent performance even in more challenging editing scenarios.
@@
Similarly, consistent trends are observed across other editing types, Edit3 (color changes) and Edit4 (object material changes) as shown in Tables 10 - 13, further reinforcing our conclusions. Regarding the FiVE-Acc metric, color changes ( Edit3) achieve the highest editing success rate among all types, with VideoGrain reaching 86% and Wan-Edit at 63%. In contrast, object material changes ( Edit4) are significantly more challenging, with AnyV2V achieving the highest success rate at 43%, followed by Pyramid-Edit at 36%. This difficulty arises because object material changes often require modifying mid- and low-frequency noise, making training-free methods highly sensitive to parameters, which leads to lower success rates.
@@
Tables 14 - 17 compare the results of Edit5 (object addition) and Edit6 (object removal). For object addition ( Edit5), VideoGrain and Wan-Edit achieve the highest scores in background preservation and motion fidelity, while TokenFlow performs best in text-vision alignment and IQA. In terms of FiVE-Acc, the RF-based methods Pyramid-Edit and Wan-Edit achieve success rates of 83% and 72%, respectively, whereas the highest-performing diffusion-based method, DMT, reaches only 61%, highlighting the advantage of RF-based approaches in this task. For Edit6 (object removal), nearly all methods perform the worst among all editing types, with FiVE-Acc scores dropping below 20%, indicating that object removal is one of the most challenging fine-grained editing tasks. This difficulty arises because removing an object requires precisely inpainting the occluded background while maintaining temporal coherence, which is particularly challenging for existing editing models. The visualizations in Fig. 4 of the main paper further confirm these findings.
@@ F More Qualitative Results and Analysis
We present the editing results across various editing types and comparison methods in Figs. 12 - 14 and Fig. 15. Figs. 12 - 14 compare the results of diffusion-based and RF-based editing methods across six editing types. These comparisons highlight the strengths and weaknesses of each method while also demonstrating the superiority of VideoGrain and our Wan-Edit. The videos shown in Fig. 15 are generated examples and represent a particularly challenging editing case. Editing becomes difficult when the object occupies a significant portion of the video frame, as it requires modifying low-frequency noise while maintaining spatial and temporal consistency. This results in failure even for the best-performing Wan-Edit. To better showcase the dynamic consistency in video editing, more video demos are available on the anonymous website: https://sites.google.com/view/five-benchmark.